# Anomalous Ionization in the Central Molecular Zone by sub-GeV Dark Matter

### v1

@@ Appendix B Ionization rate for fixed cross sections
Figure 4: Expected ionization rate in the CMZ from annihilation of sub-GeV DM particles into $e^{+}e^{-}$, as obtained with the numerical calculation for different DM distributions. The band illustrates the range of observed values of ionization rate in the CMZ.
Appendix C X and gamma ray emissions from annihilation of sub-GeV DM
@@
In the left panel of Fig. 5, we compare the predicted X-ray emission (coming mostly from inverse Compton scattering of the DM-induced $e^{\pm}$ population with CMB and other interstellar radiation fields) from the annihilation of $1$ and $10$ MeV DM for different profiles and for $\langle\sigma v\rangle=10^{-35}$ cm3/s, with the most constraining data set of the XMM-Newton measurements derived in Ref. Foster et al. (2021), the Ring 3 De la Torre Luque et al. (2023d, c). The emission from the $e^{\pm}$ particles injected by $10$ MeV DM peak at a few keV, while the peak emission for the $1$ MeV DM products peaks below the keV Cirelli et al. (2023b); De la Torre Luque et al. (2023d). Since the infrared, optical and UV radiation fields are more concentrated around the GC, the inverse-Compton emission from the scattering with these fields is more intense for the cuspier DM profiles. As a result, annihilation of sub-GeV DM can explain the observed ionization rate at the CMZ without being in conflict with X-ray emissions.
@@
Similarly, in the right panel of this figure we show the predicted MeV emission, dominated by bremsstrahlung, which is calculated with the Hermes code Dundovic et al. (2021), and compare it to SPI (INTEGRAL) measurements of the diffuse gamma-ray flux.
@@
Figure 5: Left panel: X-ray emission associated to the $e^{\pm}$ population from DM compared to XMM-Newton data in the most constraining ring (Ring 3) Foster et al. (2021). Right panel: Bremsstrahlung emission associated with the $e^{\pm}$ population injected by DM compared to the diffuse MeV gamma-ray data from SPI (INTEGRAL) Bouchet et al. (2010); Siegert et al. (2016) in the $|b|<15^{\circ}$, $|l|<30^{\circ}$ region. The gamma-ray flux at energies above the DM mass is due to reacceleration, that boosts the $e^{\pm}$ population to higher energies De la Torre Luque et al. (2023d). In both cases, these estimations are made for a $1$ and $10$ MeV DM particle for $\langle\sigma v\rangle=10^{-32}$ cm3/s, for different DM profiles

### v3

@@ Appendix C Ionization rate for fixed cross-sections
Figure 5: Expected ionization rate in the CMZ from annihilation of sub-GeV DM particles into $e^{+}e^{-}$, as obtained with the numerical calculation and for different DM distributions. The band illustrates the range of observed values of ionization rate in the CMZ. In particular, we show the average ionization rate within the CMZ for $\langle\sigma v\rangle=10^{-35}$ cm3/s. The different colors refer to the DM distributions indicated in the legend.
Appendix D X and gamma ray emissions from annihilation of sub-GeV DM
@@
In the left panel of Fig. 6, we compare the predicted X-ray emission (which comes mostly from inverse Compton scattering of the DM-induced $e^{\pm}$ population with CMB and other interstellar radiation fields) from the annihilation of $1$ and $10$ MeV DM for different profiles and for $\langle\sigma v\rangle=10^{-35}$ cm3/s, with the most constraining data set of the XMM-Newton measurements derived in Ref. Foster et al. (2021), the Ring 3 De la Torre Luque et al. (2023d, c). The emission from the $e^{\pm}$ particles injected by $10$ MeV DM peak at a few keV, while the peak emission for the $1$ MeV DM products peaks below the keV scale Cirelli et al. (2023b); De la Torre Luque et al. (2023d). Since the infrared, optical and UV radiation fields are more concentrated around the GC, the inverse-Compton emission from the scattering with these fields is more intense for the cuspier DM profiles. As a result, annihilation of sub-GeV DM can explain the observed ionization rate at the CMZ without being in conflict with X-ray emissions. We remark here that the strongest constraints so far on the annihilation rate of sub-GeV DM come from XMM-Newton measurements Cirelli et al. (2023b); De la Torre Luque et al. (2023d) of the X-ray Galactic diffuse emission, impact of the $e^{\pm}$ injection from DM annihilation on the CMB anisotropies Slatyer (2016); Lopez-Honorez et al. (2013), and the longitude profile of the $511$ keV line De la Torre Luque et al. (2023a). However, there are systematic uncertainties in the derivation of these limits that make them be less robust than the CMB limits.
@@
Similarly, in the right panel of this figure we show the predicted MeV emission, dominated by bremsstrahlung, which is calculated with the Hermes code Dundovic et al. (2021), and compare it to SPI (INTEGRAL) measurements of the diffuse $\gamma$-ray flux. We warn the reader that here we are comparing with the intensity averaged over a large region of the Galaxy. Measurements focused on the CMZ are expected to lead to a significantly higher intensity. According to our results, this does not appear as an issue for the DM hypothesis.
@@
Figure 6: Left panel: X-ray emission associated to the $e^{\pm}$ population from DM compared to XMM-Newton data in the most constraining ring (Ring 3) Foster et al. (2021). Right panel: Bremsstrahlung emission associated with the $e^{\pm}$ population injected by DM compared to the diffuse MeV $\gamma$-ray data from SPI (INTEGRAL) Bouchet et al. (2010); Siegert et al. (2016) in the $|b|<15^{\circ}$, $|l|<30^{\circ}$ region. The $\gamma$-ray flux at energies above the DM mass is due to reacceleration, that boosts the $e^{\pm}$ population to higher energies De la Torre Luque et al. (2023d). In both cases, these estimations are made for a $1$ and $10$ MeV DM particle for $\langle\sigma v\rangle=10^{-32}$ cm3/s, for different DM profiles